### v1

@@ Abstract
Routing, switching, and the interconnect fabric are essential components in implementing large-scale neuromorphic computing architectures. While this fabric plays only a supporting role in the process of computing, for large AI workloads, this fabric ultimately determines the overall system’s performance such as energy consumption and speed. In this paper we offer a potential solution to address this bottleneck by addressing two fundamental questions: (a) What computing paradigms are inherent in existing routing, switching, and interconnect systems, and how can they be used to implement a processing-in-Interconnect ($\pi^{2}$) computing paradigm ? and (b) Leveraging both the current and the future trends in interconnect technology, how will the performance of a $\pi^{2}$ system scale compared to other neuromorphic architectures? To address the first question, we demonstrate that all operations required for typical AI workloads can be mapped onto delays, causality, time-outs, packet-drop, and broadcast operations, all of which are already implemented in current packet-switching and packet-routing hardware. We show that existing buffering and traffic-shaping embedded algorithms can be leveraged to implement neuron models and synaptic operations. Additionally, a knowledge distillation framework can be used to train and cross-map well-established neural network topologies onto $\pi^{2}$ architectures without any degradation in the generalization performance. To address the second question, we use analytical modeling to show that, unlike other neuromorphic computing platforms, the energy scaling of the $\pi^{2}$ architecture improves with respect to the bandwidth, latency, and energy efficiency of the interconnects. We predict that by leveraging trends in interconnect technology, a $\pi^{2}$ architecture can be more easily scaled to execute brain-scale AI inference workloads with power consumption levels in the range of hundreds of watts.
@@
Routing, Switching, Interconnects, neuromorphic computing, AI, Network simulators
@@ 1 Introduction
While advances in artificial intelligence and neuromorphic hardware accelerators are being driven by faster and more energy-efficient computing platforms [1, 2, 3, 4, 5, 6, 7] and integrated memory technology [8, 9], an often underappreciated trend is the rapid evolution of switching and interconnect technology. Modern Ethernet switches now deliver aggregate data rates exceeding 51.2 Tera-bits-per-second with energy-efficiencies better than 10pJ/bit [10], with projections forecasting performance reaching 4.096 Peta-bits-per-second at 1pJ/bit by 2034 [11]. These high-throughput routing fabrics, whether integrated on heterogeneous packaging substrates [12] or deployed as standalone switching units [13], are essential for scaling AI and neuromorphic systems to brain-scale workloads and beyond [14]. Despite their critical role, interconnects increasingly pose a bottleneck to the overall system performance and energy-efficiency. They introduce latency [15], jitter [16], and congestion-induced packet loss, resulting in unpredictable behavior and degraded system reliability [17]. More significantly, interconnects dominate the energy budget in large-scale AI systems, as data movement consumes substantially more energy than computation and memory read/write itself. For instance, transmitting data across a two-dimensional integrated circuit can require up to 80 times more energy than executing an equivalent computation [18].
@@
Several computational paradigms have been proposed in the literature to address interconnect bottlenecks, as illustrated in Fig. 1(A)-(C). Most of these methods physically co-locate compute and memory units, aiming to alleviate bandwidth constraints and to improve energy efficiency. For example, Fig.1(B) illustrates a compute-in-memory (CIM) architecture, where matrix-vector operations are performed directly within the memory arrays, significantly reducing data movement. While the analog and digital variants of the CIM paradigm can be implemented at wafer-scale [19], scaling limitations and the heterogeneous nature of AI workloads requires re-partitioning of the compute, CIM and memory functions into distributed units. The result is a neuromorphic architecture which is shown in Fig. 1(C) which comprises of distributed cores with tightly integrated memory and compute units interconnected via a communication fabric. Even in this distributed architecture, the communication fabric acts as the dominant performance and energy bottleneck [14]. For instance, simulating one billion neurons on a large-scale system using TrueNorth processors is projected to consume approximately 4kW of power, with only 300W attributed to computation and the remainder consumed by high-performance networking switches and power supplies [20]. At brain scale, TrueNorth systems require several hundred kilowatts solely for communication and configuration infrastructure [12]. Similarly, simulating one billion neurons—approximately 1% of the human brain—on a SpiNNaker platform using 50,000 chipsets (each consuming between 0.5 and 0.76W) results in a total system power consumption exceeding 25kW. In SpiNNaker, the power dissipation due to the synaptic operations scales quadratically with the number of neurons and correlates well with inter-core communication activity. Whereas, the power dissipated at the neurons scales only linearly [21]. To quantify this bottleneck due to communication infrastructure, we define an effective energy utilization metric, $\eta$, as the ratio of energy used for actual computation to the total system energy consumed (including both compute and communication). This metric is shown in Fig. 1(F) for the architectures shown in Fig. 1(A)-(C). Details of calculating $\eta$ based on published data can be found in the Methods section 4.1. For conventional neuromorphic architectures, even though $\eta$ is projected to increase (due to improved efficiency of the communication links), as shown in Fig. 1(D), $\eta$ will still be much lower that 1.
@@
In this paper, we propose a processing-in-interconnect ($\pi^{2}$) compute architecture, where rather than considering interconnects as an overhead, the interconnect primitives are repurposed and utilized for computing. In this regard, the $\pi^{2}$ paradigm is neuromorphic in its essence, since biological brains are also hypothesized to exploit axonal and dendritic delays for computing [18, 22]. However, in $\pi^{2}$ architecture, our goal is to leverage the interconnect primitives that already exists or is efficiently implemented in current routing, switching and interconnect hardware. This will obviate the need for developing new custom hardware and at the same time support scaling to larger workloads. These interconnect primitives are summarized in Fig. 1F and their equivalence to neural network operations is described in the Methods section  4.2. For instance, the delay operation is shown to be equivalent to a MULTIPLY operation and a memory operation, the sorting operation is shown to be equivalent to ADD operation and event time-out/drop-out is shown to be equivalent to the non-linear activation. Since we will be considering only packet switching technologies for implementing the interconnect fabric, events and data packets will be used interchangeably. In this regard, $\pi^{2}$ architecture would also be able to leverage the event routing infrastructure in existing address-event-representation (AER) neuromorphic infrastructure [23]. Since the computing and the memory functions are integrated with the interconnect functions as shown in Fig. 1E, the energy-utilization metric $\eta$ for the $\pi^{2}$ architecture is not constrained by the interconnect bottlenecks. Increasing the interconnect bandwidth effectively scales the computational bandwidth, and since interconnects also serve as memory elements, they alleviate the memory-bandwidth bottlenecks in conventional architectures [24, 25] (Supplementary section  1). As a result, the effective energy-utilization metric $\eta$ for the $\pi^{2}$ architecture is close to 1, as shown in Fig. 1F. The key contribution of this work is to show how the $\pi^{2}$ architecture can be implemented and scaled using packet-switching networks and using well established network protocols and routing primitives; and how the $\pi^{2}$ architecture can be trained to execute typical AI tasks.
@@ 1.1 Credit-based traffic shaping acts as $\pi^{2}$ Neuron
Figure 2: A) Illustration of traditional Credit-Based Shaper (CBS) operation where each traffic class (e.g., $G_{1}$, $G_{2}$) is associated with an individual queue and a CBS module. Events are dropped if the queue overflows. Transmission occurs only when the associated credit is non-negative and the channel is free. In ii), Credit dynamics for two traffic classes: $G_{1}$ (red) and $G_{2}$ (blue) is shown. AE and TE represent the arrival and transmission of events, respectively. Initially, the channel is occupied (black region). So the arriving event of class $G_{1}$ has to be queued. Accordingly, its credit starts increasing. Once the channel is free, the packet corresponding to $G_{1}$ gets transmitted, and its credit decreases with a slope. Similar dynamics is plotted for $G_{2}$. B) i) The $\pi^{2}_{K}$ neuron integrates the first K temporal input events and generates a single output event based on a time-to-first-spike encoding scheme. ii) As input events arrive at times $T_{1}$ to $T_{4}$, the membrane potential increases with a slope proportional to the number of inputs received. After receiving $K$ inputs, the slope remains constant. When the membrane potential crosses a predefined threshold $M$, the neuron emits an output event at $T$. The dynamics of a $\pi^{2}_{3}$ neuron is plotted here. iii) $\pi^{2}$-neuron - The figure illustrates the operation of the proposed CBS protocol, which aligns the credit-based traffic shaping (CBS) protocol with the dynamics of the $\pi^{2}_{K}$ neuron model. The modified credit dynamics of a $\pi^{2}_{2}$ neuron for the traffic classes $G_{1},G_{2}$ is highlighted here. The bottom panels contrast traditional credit-based dynamics with the proposed $\pi^{2}_{K}$ dynamics. C) Traditional Asynchronous Traffic Shaping (ATS) mechanism: Incoming events are assigned to $2^{P}$ shaped queues based on their Priority Code Point (PCP) values. The ATS shaper delays each packet until its Transmission Eligibility Time (TET), calculated usually as a function of the traffic characteristics, is reached. Once eligible, events are forwarded to a shared queue (which is controlled by the modified CBS protocol to emulate the $\pi^{2}_{K}$ neuron dynamics) in a time-sorted order for transmission. The ATS protocol is re-envisioned to model the $\pi^{2}$ synapse. For the $i^{th}$ event, the $TET_{i}$ is computed as $T_{i}+W_{i}$, where $W_{i}$ is the synaptic delay (PCP code) associated with it. D) Illustration of a fully connected (n x n) $\pi^{2}$-NN architecture implemented using $\pi^{2}_{K}$ neurons, where synaptic weights are encoded as interconnect delays ($W_{1}$, $W_{2}$, $W_{3}$, …).
@@
Ethernet switches are designed to handle a diverse mix of traffic types, ranging from time-sensitive audio/video streams, to control messages and best-effort data. The switches also incorporate built-in mechanisms to support differentiated treatment of these different types of traffic. To manage this heterogeneity, switches implement multiple output queues at each port, with each queue corresponding to a specific traffic class or priority level. The flow of data through these queues is regulated by traffic shaping protocols, which control the timing and rate of packet transmission. A credit-Based Shaping (CBS) protocol regulates transmission by associating a dynamic credit counter with each queue [26, 27] as shown in Fig. 2A. A frame from a queue is eligible for transmission only when its corresponding credit is non-negative. Credits are updated based on the state of the queue: they accumulate at a fixed (idle) rate when frames are waiting, and decrease (send slope) during active transmission. This behavior is analogous to the integration and leakage dynamics in spiking neurons, where the membrane potential builds up before an event is generated and then resets or decays after the event.
@@
We propose a modification of the CBS algorithm to implement a $\pi^{2}_{K}$ neuron model. Specifically, the $\pi^{2}_{K}$ neuron model uses a variant of the time-domain margin-propagation (TEMP) algorithm [23] and in the Methods section 4.2 we describe their equivalence. The $\pi^{2}_{K}$ neuron generates an event at time $T$, that is computed as follows;
@@
Let $\{t_{1},t_{2},\ldots,t_{d}\}$ denote the set of input event arrival times, where $d$ is the total number of inputs received by the neuron. The terms $t_{[1]},t_{[2]},\ldots,t_{[K]}$ represent the first $K$ earliest arrival times among the inputs. Thus, the neuron computes the mean of the earliest $K$ inputs and adds a constant offset $M/K$, resulting in the output time $T$. The constant $M/K$ ensures causality such that $T\geq t_{[K]}$. In a CBS-based model, the credit accumulation mirrors the neuron’s membrane potential dynamics, where the number of incoming events modulates the accumulation rate (slope), represented here by the number of frames buffered in a shared queue. When the accumulated credit crosses a predefined threshold (M), the queue gate is enabled, and the buffered events are transmitted, analogous to a neuron’s firing event. Also, every shared queue is sized to accommodate only $K$ frames/events. The remaining incoming events are automatically dropped as portrayed in Fig. 2A. If the threshold is not reached before a pre-defined time out value $T_{out}$, then all the events in the shared queue are dropped, and the credits are reset. The dynamics of the $\pi^{2}_{K}$ neuron is depicted in Fig. 2B.
@@ 1.2 Asynchronous traffic shaping act as $\pi^{2}$ synapses
The Asynchronous Traffic Shaping (ATS) protocol is another traffic shaping algorithm defined in the IEEE 802.1Qcr [28], designed for regulating the transmission of frames in Ethernet switches—particularly in time-sensitive networks. Unlike time-aware shapers that rely on global time synchronization, ATS operates asynchronously by assigning a Transmission Eligibility Time (TET) to each frame buffered in the output queues [28] (Fig. 2C). The TET determines when a frame becomes eligible for transmission, based on local parameters such as frame arrival time, burst, and frame-size. In this regard, the ATS protocol facilitates precise temporal coordination without relying on centralized synchronization. Each synaptic delay is quantized and encoded into a $p$-bit Priority Code Point (PCP) [27]. The PCP field, defined in the IEEE 802.1Q VLAN tag (usually a 3-bit field), enables traffic differentiation by assigning each frame a priority level. Ethernet switches utilize this field to select the appropriate queue, apply shaping policies, and prioritize time-sensitive traffic. To enable delay-based event communication within the $\pi^{2}$ network, the modified ATS protocol supports three core neuromorphic operations:
@@
We extend the Asynchronous Traffic Shaping (ATS) algorithm to compute Transmission Eligibility Time (TET) values based on the synaptic delay parameters in a $\pi^{2}$ neural network. An example of a fully connected ($N\times N$) instantiation of a $\pi^{2}$ neural network is shown in Fig.2D. Let $T_{i}$ denote the event generation time of the $i^{\text{th}}$ presynaptic (Tx) neuron, and let $W_{ij}$ represent the synaptic delay from Tx neuron $i$ to postsynaptic (Rx) neuron $j$. The corresponding arrival time $\tau_{ij}$ of the event from neuron $i$ at neuron $j$ is given by:
@@
In any processor architecture, delays are an unavoidable consequence of architectural features such as memory access latency, interconnect traversal time, and scheduling overhead. Rather than treating these delays as inefficiencies, the $\pi^{2}$ framework reinterprets them as functional components of synaptic processing. Fig. 3A illustrates a hardware-aligned approach on how these intrinsic delays are exploited through a combination of hierarchical routing and modified traffic shaping protocols. The synaptic delay can be decomposed as a sum of the memory access delay ($d_{v}$), determined by the depth of the routing structure, and a queuing delay ($W_{ij}^{\prime}$), determined by the ATS protocol. The delayed event is then processed by the modified CBS protocol, which emulates the integration and thresholding behavior of the $\pi^{2}_{K}$ neuron. Once an output event is generated, it is fed back through the same process of hierarchical routing and traffic shaping pipeline. The routing table determines the overall network topology (feedforward, recurrent or hybrid) and the recursive event-driven computation with memory-access delay and delay-based scheduling proceeds till the completion of the inference task. For instance in a feedforward network, the total recursion time is determined by the depth of the network.

### v3

@@ Abstract
Routing, switching, and the interconnect fabric are essential components in implementing large-scale neuromorphic computing architectures. While this fabric plays only a supporting role in the process of computing, for large AI workloads, this fabric ultimately determines the overall system’s performance, such as energy consumption and speed. In this paper, we offer a potential solution to address this bottleneck by addressing two fundamental questions: (a) What computing paradigms are inherent in existing routing, switching, and interconnect systems, and how can they be used to implement a Processing-in-Interconnect ($\pi^{2}$) computing paradigm? and (b) how to train $\pi^{2}$ network on standard AI benchmarks ? To address the first question, we demonstrate that all operations required for typical AI workloads can be mapped onto delays, causality, time-outs, packet-drop, and broadcast operations, all of which are already implemented in current packet-switching and packet-routing hardware. We then show that existing buffering and traffic-shaping embedded algorithms can be minimally modified to implement $\pi^{2}$ neuron models and synaptic operations. To address the second question, we show how a knowledge distillation framework can be used to train and cross-map well-established neural network topologies onto $\pi^{2}$ architectures without any degradation in the generalization performance. Our analysis show that the effective energy utilization of a $\pi^{2}$ network is significantly higher than that of other neuromorphic computing platforms, as a result we believe that the $\pi^{2}$ paradigm offers a more scalable architectural path toward achieving brain-scale AI inference.
@@
Routing, Switching, Interconnects, neuromorphic computing, artificial intelligence, Network simulators
@@ 1 Introduction
While advances in artificial intelligence and neuromorphic hardware accelerators are being driven by faster and more energy-efficient computing platforms [h100, tpu, tpuv6, nvidiacpu, nvidiacpu1, npu, amdfpga, andre] and integrated memory technology [tsun, hbm_pim], an often underappreciated trend is the rapid evolution of switching and interconnect technology. Modern Ethernet switches now deliver aggregate data rates exceeding 51.2 Tera-bits-per-second with energy-efficiencies better than 10pJ/bit [marvel_tera], with projections forecasting performance reaching 4.096 Peta-bits-per-second at 1pJ/bit by 2034 [cisco]. These high-throughput routing fabrics, whether integrated on heterogeneous packaging substrates [truenorth_large_scale] or deployed as standalone switching units [cisco_one_family], are essential for scaling AI and neuromorphic systems to brain-scale workloads and beyond [large_scale_neuro]. Despite their critical role, interconnects increasingly pose a bottleneck to the overall system performance and energy efficiency. They introduce latency [latency], jitter [jitter], and congestion-induced packet loss, resulting in unpredictable behavior and degraded system reliability [photonic]. More significantly, interconnects dominate the energy budget in large-scale AI systems, as data movement consumes substantially more energy than computation and memory read/write itself. For instance, transmitting data across a two-dimensional integrated circuit can require up to 80 times more energy than executing an equivalent computation [dendrocentric].
@@
Several computational paradigms have been proposed in the literature to address interconnect bottlenecks, as illustrated in Fig. 1(A)-(C). Most of these methods physically co-locate compute and memory units, aiming to alleviate bandwidth constraints and to improve energy efficiency. For example, Fig.1(B) illustrates a compute-in-memory (CIM) architecture, where matrix-vector operations are performed directly within the memory arrays, significantly reducing data movement [jason, tsun]. While the analog and digital variants of the CIM paradigm can be implemented at wafer-scale [CEREBRAS], scaling limitations and the heterogeneous nature of AI workloads require re-partitioning of the compute, CIM, and memory functions into distributed units. The result is a neuromorphic architecture, which is shown in Fig. 1(C), which comprises of distributed cores with tightly integrated memory and compute units interconnected via a communication fabric [brainscales]. Even in this distributed architecture, the communication fabric acts as the dominant performance and energy bottleneck [large_scale_neuro]. For instance, simulating one billion neurons on a large-scale system using TrueNorth processors is projected to consume approximately 4kW of power, with only 300W attributed to computation and the remainder consumed by high-performance networking switches and power supplies [truenorth3]. At brain scale, TrueNorth systems require several hundred kilowatts solely for communication and configuration infrastructure [truenorth_large_scale]. Similarly, simulating one billion neurons—approximately 1% of the human brain—on a SpiNNaker platform using 50,000 chipsets (each consuming between 0.5 and 0.76W) results in a total system power consumption exceeding 25kW. In SpiNNaker, the power dissipation due to the synaptic operations scales quadratically with the number of neurons and correlates well with inter-core communication activity. Whereas, the power dissipated at the neuron scales only linearly [spinnaker]. To quantify this bottleneck due to communication infrastructure, we define an effective energy utilization metric, $\eta$, as the ratio of energy used for actual computation to the total system energy consumed (including both compute and communication). This metric is shown in Fig. 1(F) for the architectures shown in Fig. 1(A)-(C). Details of calculating $\eta$ based on published data can be found in the Methods section 4.1. For conventional neuromorphic architectures, even though $\eta$ is projected to increase (due to improved efficiency of the communication links), as shown in Fig. 1(D), $\eta$ will still be much lower than 1.
@@
In this paper, we propose a processing-in-interconnect ($\pi^{2}$) compute architecture, where rather than considering interconnects as an overhead, the interconnect primitives are repurposed and utilized for computing. In this regard, the $\pi^{2}$ paradigm is neuromorphic in its essence, since biological brains are also hypothesized to exploit axonal and dendritic delays for computing [dendrocentric, axonal]. However, in $\pi^{2}$ architecture, our goal is to leverage the interconnect primitives that already exist or are efficiently implemented in current routing, switching, and interconnect hardware. These interconnect primitives are summarized in Fig. 1F and their equivalence to neural network operations is described in the Methods section  4.2. For instance, the delay operation is shown to be equivalent to a MULTIPLY operation and a memory operation, the sorting operation is shown to be equivalent to an ADD operation, and event time-out/drop-out is shown to be equivalent to the non-linear activation. Since we will be considering only packet switching technologies for implementing the interconnect fabric, events and data packets will be used interchangeably. In this regard, $\pi^{2}$ architecture would also be able to leverage the event routing infrastructure in existing address-event-representation (AER) neuromorphic infrastructure [temp, gert, gert2]. Since the computing and the memory functions are integrated with the interconnect functions as shown in Fig. 1E, the energy-utilization metric $\eta$ for the $\pi^{2}$ architecture is not constrained by the interconnect bottlenecks. Increasing the interconnect bandwidth effectively scales the computational bandwidth, and since interconnects also serve as memory elements, they alleviate the memory-bandwidth bottlenecks in conventional architectures [cim1, cim2] (Supplementary section  1 and Fig. S1). As a result, the effective energy-utilization metric $\eta$ for the $\pi^{2}$ architecture is close to 1, as shown in Fig. 1F. The key contribution of this work is to show how the $\pi^{2}$ architecture can be implemented and scaled using packet-switching networks and using well-established network protocols and routing primitives with minimal modifications, and how the $\pi^{2}$ architecture can be trained to execute typical AI tasks.
@@ 1.1 Credit-based traffic shaping acts as $\pi^{2}$ Neuron
Figure 2: A) Illustration of traditional Credit-Based Shaper (CBS) operation where each traffic class (e.g., $G_{1}$, $G_{2}$) is associated with an individual queue and a CBS module. Events are dropped if the queue overflows. Transmission occurs only when the associated credit is non-negative and the channel is free. In ii), Credit dynamics for two traffic classes: $G_{1}$ (red) and $G_{2}$ (blue) are shown. AE and TE represent the arrival and transmission of events, respectively. Initially, the channel is occupied (black region). So the arriving event of class $G_{1}$ has to be queued. Accordingly, its credit starts increasing. Once the channel is free, the packet corresponding to $G_{1}$ gets transmitted, and its credit decreases with a slope. Similar dynamics is plotted for $G_{2}$. B) i) The $\pi^{2}_{K}$ neuron integrates the first K temporal input events and generates a single output event based on a time-to-first-spike encoding scheme. ii) As input events arrive at times $T_{1}$ to $T_{4}$, the membrane potential increases with a slope proportional to the number of inputs received. After receiving $K$ inputs, the slope remains constant. When the membrane potential crosses a predefined threshold $M$, the neuron emits an output event at $T$. The dynamics of a $\pi^{2}_{3}$ neuron is plotted here. iii) $\pi^{2}$-neuron - The figure illustrates the operation of the proposed CBS protocol, which aligns the credit-based traffic shaping (CBS) protocol with the dynamics of the $\pi^{2}_{K}$ neuron model. The modified credit dynamics of a $\pi^{2}_{2}$ neuron for the traffic classes $G_{1},G_{2}$ is highlighted here. The bottom panels contrast traditional credit-based dynamics with the proposed $\pi^{2}_{K}$ dynamics. C) Traditional Asynchronous Traffic Shaping (ATS) mechanism: Incoming events are assigned to $2^{P}$ shaped queues based on their Priority Code Point (PCP) values. The ATS shaper delays each packet until its Transmission Eligibility Time (TET), calculated usually as a function of the traffic characteristics, is reached. Once eligible, events are forwarded to a shared queue (which is controlled by the modified CBS protocol to emulate the $\pi^{2}_{K}$ neuron dynamics) in a time-sorted order for transmission. The ATS protocol is re-envisioned to model the $\pi^{2}$ synapse. For the $i^{th}$ event, the $TET_{i}$ is computed as $T_{i}+W_{i}$, where $W_{i}$ is the synaptic delay (PCP code) associated with it. D) Illustration of a fully connected (n x n) $\pi^{2}$-NN architecture implemented using $\pi^{2}_{K}$ neurons, where synaptic weights are encoded as interconnect delays ($W_{1}$, $W_{2}$, $W_{3}$, …).
@@
Ethernet switches are designed to handle a diverse mix of traffic types, ranging from time-sensitive audio/video streams to control messages and best-effort data. The switches also incorporate built-in mechanisms to support differentiated treatment of these different types of traffic. To manage this heterogeneity, switches implement multiple output queues at each port, with each queue corresponding to a specific traffic class or priority level. The flow of data through these queues is regulated by traffic shaping protocols, which control the timing and rate of packet transmission. A credit-based Shaping (CBS) protocol defined in the IEEE 802.1Qav standard regulates transmission by associating a dynamic credit counter with each queue [cbs, nesting] as shown in Fig. 2A. A frame from a queue is eligible for transmission only when its corresponding credit is non-negative. Credits are updated based on the state of the queue: they accumulate at a fixed (idle) rate when frames are waiting, and decrease (send slope) during active transmission. This behavior is analogous to the integration and leakage dynamics in spiking neurons, where the membrane potential builds up before an event is generated and then resets or decays after the event.
@@
We propose a modification of the CBS algorithm to implement a $\pi^{2}_{K}$ neuron model. Specifically, the $\pi^{2}_{K}$ neuron model uses a variant of the time-domain margin-propagation (TEMP) algorithm [temp] and in the Methods section 4.2 we describe their equivalence. The $\pi^{2}_{K}$ neuron generates an event at time $T$, that is computed as follows;
@@
Let $\{t_{1},t_{2},\ldots,t_{d}\}$ denote the set of input event arrival times, where $d$ is the total number of inputs received by the neuron. The terms $t_{[1]},t_{[2]},\ldots,t_{[K]}$ represent the first $K$ earliest arrival times among the inputs. Thus, the neuron computes the mean of the earliest $K$ inputs and adds a constant offset $M/K$, resulting in the output time $T$. The constant $M/K$ ensures causality such that $T\geq t_{[K]}$. In a CBS-based model, credit accumulation mirrors the neuron’s membrane potential dynamics, with the number of incoming events modulating the accumulation rate (slope), which is represented here by the number of frames buffered in a shared queue. When the accumulated credit crosses a predefined threshold (M), the queue gate is enabled, and the buffered events are transmitted, analogous to a neuron’s firing event. Also, every shared queue is sized to accommodate only $K$ frames/events. The remaining incoming events are automatically dropped as portrayed in Fig. 2A. If the threshold is not reached before a pre-defined time-out value $T_{out}$, then all the events in the shared queue are dropped, and the credits are reset. The dynamics of the $\pi^{2}_{K}$ neuron is depicted in Fig. 2B. The $\pi^{2}$ computation requires modest microarchitectural extensions to the CBS implementation—specifically programmable credit thresholds (an extension discussed in [cbs_threshold] ) and queue length-dependent credit accumulation—which extends the internal credit-update logic (implemented using a $\log_{2}(K)$-bit counter shown in Fig. S4 in the Supplementary section 3.1). Additional details of the modifications required in the conventional CBS algorithm to emulate the $\pi^{2}$ dynamics are presented in the supplementary section 3.1(Fig. S4 and Table S1). Note that these changes are only applied to the microarchitecture of the CBS shaper; the Ethernet frame formats are still compliant with the IEEE standard.
@@ 1.2 Asynchronous traffic shaping act as $\pi^{2}$ synapses
The Asynchronous Traffic Shaping (ATS) protocol is another traffic shaping algorithm defined in the IEEE 802.1Qcr [tsnintro], designed for regulating the transmission of frames in Ethernet switches—particularly in time-sensitive networks. Unlike time-aware shapers that rely on global time synchronization, ATS operates asynchronously by assigning a Transmission Eligibility Time (TET) to each frame buffered in the output queues [tsnintro] (Fig. 2C). The TET determines when a frame becomes eligible for transmission, based on local parameters such as frame arrival time, burst, and frame size. In this regard, the ATS protocol facilitates precise temporal coordination without relying on centralized synchronization. Each synaptic delay is quantized and encoded into a $p$-bit Priority Code Point (PCP) [nesting]. The PCP field, defined in the IEEE 802.1Q VLAN tag (usually a 3-bit field), enables traffic differentiation by assigning each frame a priority level. Ethernet switches utilize this field to select the appropriate queue, apply shaping policies, and prioritize time-sensitive traffic. To enable delay-based event communication within the $\pi^{2}$ network, the modified ATS protocol supports three core neuromorphic operations:
@@
We extend the Asynchronous Traffic Shaping (ATS) algorithm to compute Transmission Eligibility Time (TET) values based on the synaptic delay parameters in a $\pi^{2}$ neural network. The modification of the ATS protocol (IEEE 802.1Qcr) to support $\pi^{2}$ computation is presented in the supplementary section 3.1 (Algorithm 1). Note that the modification does not add any new variables or extra computation; it simply repurposes a precomputed value to introduce delays. An example of a fully connected ($N\times N$) instantiation of a $\pi^{2}$ neural network is shown in Fig.2D. Let $T_{i}$ denote the event generation time of the $i^{\text{th}}$ presynaptic (Tx) neuron, and let $W_{ij}$ represent the synaptic delay from Tx neuron $i$ to postsynaptic (Rx) neuron $j$. The corresponding arrival time $\tau_{ij}$ of the event from neuron $i$ at neuron $j$ is given by:
@@
In any processor architecture, delays are an unavoidable consequence of architectural features such as memory access latency, interconnect traversal time, and scheduling overhead. Rather than treating these delays as inefficiencies, the $\pi^{2}$ framework reinterprets them as functional components of synaptic processing. Fig. 3A illustrates a hardware-aligned approach on how these intrinsic delays are exploited through a combination of hierarchical routing and modified traffic shaping protocols. The synaptic delay can be decomposed as a sum of the memory access delay ($d_{v}$), determined by the depth of the routing structure, and a queuing delay ($W_{ij}^{\prime}$), determined by the ATS protocol. The delayed event is then processed by the modified CBS protocol, which emulates the integration and thresholding behavior of the $\pi^{2}_{K}$ neuron. Once an output event is generated, it is fed back through the same process of hierarchical routing and traffic shaping pipeline. The routing table determines the overall network topology (feedforward, recurrent, or hybrid) and the recursive event-driven computation with memory-access delay and delay-based scheduling proceeds till the completion of the inference task. For instance, in a feedforward network, the total recursion time is determined by the depth of the network.